One mole of O2(g), with CV,m=2.5R,  is expanded adiabatically from 301 K and 4.00 bar to 1.60 bar against a constant external pressure equal to the final pressure. Calculate q, w, ΔU, ΔH, ΔS and ΔSsurr.
Enter your answers in the specified units with three or more significant figures . Do not include units as part of your answer.

   q = ____________ J
   w = ____________ J
ΔU = ____________ J
ΔH = ____________ J
ΔS = ____________ J mol-1 K-1
ΔSsurroundings = ____________ J mol-1 K-1
Answer #1

γουν constant peessue, the pe Ansit given that, that the process is adiabatic and as the process is an expansion of the gas aVOUV WE - 2.5 X8:314 ( 301 - 249:40 W = -1072-506J/ step-111 : calculate AU AU = CudT . But dT = To T = 249.40-301 = -51.6k.peessure & step-v (calculate as: Entropy_change as a function of temp. as is given by the ear- AS: 2.303 (pdog. t + 2:303 R J